# Generating variable $\hbar$ and $c$ via Fujii–Wetterich model in curved spacetimes

### v1

@@ Abstract
We revisit the Fujii–Wetterich model [Phys.Rev.D 26, 2580 (1982) and Nucl.Phys.B 302, 645 (1988)] which allows the Higgs doublet to couple with a ‘cosmon’ scalar $\chi$ of the background spacetime as $\chi^{2}\,\Phi^{2}$. Upon spontaneous breaking of the $SU(2)$ gauge symmetry, the vacuum expectation value (VEV) of the Higgs doublet is proportional to the ‘cosmon’ field $\chi$. Fujii and Wetterich employed this linkage to make particle masses dependent on $\chi$. We shall present an alternative mechanism: at a given point $x^{*}$, the prevailing Higgs VEV will be used to construct a quantum of action $\hbar_{*}$ and a speed of light $c_{*}$ in association with $\chi(x^{*})$. Specifically, each open set vicinity of a point $x^{*}$ on the manifold is equipped with a replica of the Glashow–Weinberg–Salam action operating with its own effective values of $\hbar_{*}$ and $c_{*}$, whereas particle masses induced via Higgs spontaneous symmetry breaking remain independent of $\chi(x^{*})$. Our mechanism unambiguously generates the dependencies $\hbar_{*}\propto\chi^{-1/2}(x^{*})$ and $c_{*}\propto\chi^{1/2}(x^{*})$, causing these ‘fundamental constants’ to vary along with the dynamical field $\chi$ across the manifold. Moreover, in each open set around $x^{*}$, the field $\chi(x^{*})$ determines the length and time scales for physical processes occurring in this region as follows: $l\propto\chi^{-1}(x^{*})$ and $\tau\propto\chi^{-3/2}(x^{*})$. The resulting anomalous anisotropic relation between the rate of clocks and the length of rods, $\tau^{-1}\propto l^{-3/2}$, leads to a concrete physical prediction, manifesting as a new time dilation effect, and produces a distinct set of previously unknown phenomenology.
For late-time cosmology, the variation of $c$ along the trajectory of light waves from distant supernovae towards the Earth-based observer renders the classic Lemaître redshift formula $1+z=a^{-1}$ inapplicable. Using the dependency $c_{*}\propto\chi^{1/2}(x^{*})$, we derive the new (variable-$c$) Lemaître redshift relation and apply it to analyze the Pantheon Catalog of Type Ia supernovae without invoking the dark energy hypothesis. Detailed expositions of our analysis are provided in companion papers, with the key consequences summarized herein: (1) Accounting for the Pantheon Catalog with a fit exceeding the quality of the $\Lambda$CDM model; (2) Explaining the late-time cosmic acceleration based solely on variable $c$, eliminating the need for dark energy; (3) Revitalizing Blanchard–Douspis–Rowan-Robinson–Sarkar’s CMB power spectrum analysis that bypassed dark energy [A&A 412, 35 (2003) and A&A 449, 925 (2006)]; (4) Resolving the age problem without resorting to dark energy; and (5) Resolving the $H_{0}$ tension through the evolution of galaxies on an expanding cosmic background, rather than requiring a dynamical dark energy component.
I Theoretical consideration
Motivation—In the light of several outstanding issues and tensions in cosmology, there has emerged growing interest in promoting the parameters of physical models to be variable scalar fields in spacetime. One popular example is the dynamical dark energy which could offer relief to the fine tuning and coincidence problems in late-time cosmology. The cosmological constant $\Lambda$ is equipped with a kinetic term that allows it to evolve and ‘relax’ to its (small) value in the current epoch.
@@ I Theoretical consideration
where $\omega$ is a dimensionless (BD) parameter. The BD theory has grown into a family of scalar–tensor theories, a very popular theme of research nowadays. The field $\chi^{2}$, often known a ‘dilaton’, also naturally arises in string theory and other contexts [5].
On the face value, the (Newton) gravitational ‘constant’ becomes a scalar field via its dependency on the field $\chi$, via a formal identification
$\frac{1}{16\pi}\chi^{2}\,\mathcal{R}:=\frac{c^{3}}{16\pi\hbar G}\,\mathcal{R}\ \ \ \Longrightarrow\ \ \ G\propto\chi^{-2}\,.$ (3)
With the aid of Eq. (1), this dependency results in
$m_{P}\propto\chi;\ \ \ l_{P}\propto\chi^{-1};\ \ \ \tau_{P}\propto\chi^{-1}$ (4)
In the vanilla BD theory, the matter sector coupled minimally to the gravitational sector to preserve Einstein’s Equivalence Principle [4]. However, around the same time, Utiyama and DeWitt showed that even if one starts with a minimal coupling action between matter and gravity, radiative corrections would generate non-minimal coupling between them [6]. Since the 1980s, efforts to introduce non-minimal coupling directly at tree level have been pursued independently in the works of Fujii and Wetterich, which we shall call collectively as the Fujii–Wetterich scheme. In [7] Fujii allowed the Higgs doublet $\Phi$ of the Glashow–Weinberg–Salam (GWS) model of particle physics to couple with the Ricci scalar in the form $\mathcal{R}\,\Phi^{2}$. In [8, 9, 10, 11, 12] Wetterich embedded the BD action in a broader framework that projects a dilatation symmetry, with the non-minimal coupling in the form  [^1] Wetterich referred the field $\chi$ a ‘cosmon’ [8]. In place of Eq. (5), he considered a more general form $V(\chi,\Phi)=\Phi^{4}\,w\left(\chi^{2}/\Phi^{2}\right)$ where $w$ is a function of the ratio $x:=\chi^{2}/\Phi^{2}$. If $w(x)$ adopts an affine form $\lambda-\mu^{2}\,x$, then Eq. (5) would ensue.
@@
where $\lambda\in\mathbb{R^{+}}$ and $\mu\in\mathbb{R}$ are dimensionless parameters.
@@
An essential element in the Fujii–Wetterich scheme is that, when the $SU(2)$ gauge symmetry of the Higgs doublet undergoes spontaneous breaking by way of the potential prescribed in Eq. (5), the field $\chi$ influences the vacuum expectation value (VEV) of the Higgs doublet $\Phi$, which in turn affects the mass of the fermions and the gauge vector bosons in the GWS model. Specifically, Fujii and Wetterich demonstrated that the electron mass $m$ scales proportional to $\chi$, thereby maintaining a constant ratio $m/m_{P}$. Moreover, the length scale $l$ and time scale $\tau$ of any given physical process also become dependent on $\chi$ as $l\propto\chi^{-1}$ and $\tau\propto\chi^{-1}$, in exact proportion to $l_{P}$ and $\tau_{P}$ respectively. Most importantly, the field $\chi$ affects not only the Planck time per Eq. (4) but also the rate of the clocks (viz. the revolution rate of a mechanical clock, the oscillation rate of an atomic clock, or the decay rate of an unstable quantum system) which are made of matter governed by the GWS model. Quantitatively, the Fujii–Wetterich scheme predicts that the evolution rate $\tau^{-1}$ of physical processes is proportional to $\chi$.  [^2] We find it illuminating to quote Fujii [7]: “… the time and length in the microscopic unit frame are measured in units of $m^{-1}(t)$, in agreement with the physical situation that the time scale of atomic clocks, for example, is provided by the atomic levels which are determined by the Rydberg constant $(me^{4})^{-1}$ and Wetterich [10]: “The clock provided by the Hubble expansion in the standard description is now replaced by a clock associated to the increasing value of $\chi$.
@@
It is imperative to stress that the field $\chi$ itself serves as a (dynamical) length scale that can be used for constructing $G$, a dimensionful quantity. It is then logically permissible to adopt an alternative mechanism whereby the field $\chi$, a quantity equal to $\sqrt{c^{3}/(\hbar G)}$, could also be used to derive $c$ and $\hbar$, rather than $G$. The key objective of our paper is to establish the dependencies of $\hbar$ and $c$ on $\chi$ via the matter sector. Our approach constitutes a major departure from the Fujii–Wetterich scheme: by requiring the charge and mass of elementary particles to be independent of $\chi$, we unambiguously obtain $c\propto\chi^{1/2}$ and $\hbar\propto\chi^{-1/2}$ while $G$ remains constant. Furthermore, via the time evolution of quantum states $i\hbar\frac{\partial}{\partial t}\left|\psi\right\rangle=\hat{H}\left|\psi\right\rangle$, the dependency $\hbar\propto\chi^{-1/2}$, along with $\hat{H}\propto\chi$, causes the clock rate of physical processes to scale as $\tau^{-1}\propto\chi^{3/2}$, in decisive distinction from the Fujii–Wetterich scheme (which posits $\tau^{-1}\propto\chi$).
@@
The anisotropic time scaling enabled in our approach, characterized by its anomalous $3/2$–exponent, leads to new physics with a distinct set of phenomenology andpredictions. We will discuss these significant implications in this paper and its companions [13, 14, 15, 16].
@@
History of variable $c$ and $\hbar$—Prior to the Brans-Dicke theory, Dicke also briefly considered variability in the speed of light $c$ in 1957 [17] although he did not pursue it further. A relatively little-known fact is that the concept of variable speed of light (VSL) was initially explored by Einstein in 1911, published in three obscure papers (originally in German) [18, 19, 20]  [^3] It seems that Dicke in 1957 [17] was not aware of Einstein’s 1911 VSL attempt [18].. During his quest for a generally covariant theory of gravity (which later became known as the General Relativity), Einstein emphasized that the Michelson–Morley experimental results and Lorentz symmetry are meant to hold only locally. That is to say, while $c$ is invariant with respect to local boosts, its value needs not be universal. Einstein envisioned the possibility that gravitational potential affected the clock rate, hence generating a variable $c$ in spacetime. Such a variation in $c$ could then produce a refraction effect on light ray around the Sun’s disc as a result of Huygens’s principle [18]. The later successes of General Relativity quickly overshadowed Einstein’s 1911 paper on VSL, leading him to abandon the idea, however. In the 1990s, VSL was revived independently by Moffat and by Albrecht and Magueijo to tackle issues in early-time cosmology, such as the horizon paradox [21, 22].
@@
The concept of variable quantum of action $\hbar$ has been much less explored, with a rare exception of [23]. One important empirical guidance is that, although the fine-structure constant $\alpha:=e^{2}/(\hbar c)$ is known to ‘run’ in the renormalization group (RG) flow of quantum field theory, there is no clear experimental evidence supporting $\alpha$ varying in spacetime  [^4] In [24] Webb et al. offered observational evidence of a varying $\alpha$ with respect to the redshift, but this result has been in contention; e.g., see a recent paper [25].. If $\alpha$ is to remain independent of the field $\chi$ while $c$ is allowed to vary, it would require a co-variation of $e$, $\hbar$, and $c$. One such scenario, as briefly mentioned earlier, posits $\hbar\propto\chi^{-1/2}$ and $c\propto\chi^{1/2}$, which would insulate $\alpha$ from depending on $\chi$.
A Higgs–scalon coupling model—Consider the full action
@@
in which the matter sector is specified as
@@
where all parameters $\alpha$, $f$, $\mu$, and $\lambda$ are dimensionless. The matter sector comprises a massless spinor field $\psi$ coupled with a massless $U(1)$ gauge field $A_{\mu}$, and a charge-neutral Higgs singlet $\Phi$ coupled with a scalar singlet $\chi$ on a 4-dimensional Riemannian spacetime manifold. We shall term the field $\chi$ a ‘scalon’ (a concept to be clarified later), representing a scalar degree of freedom within the gravitational sector.
It is important to remark that the adimensional action of the matter sector does not involve the (dimensionful) quantum of action $\hbar$ and speed of light $c$ at its outset. The length dimensions of $\{\psi,\,A_{\mu},\,\Phi$} are $\{-3/2,-1,-1\}$ respectively. The gamma matrices satisfy $\{\gamma^{\mu}(x),\gamma^{\nu}(x)\}=2\,g^{\mu\nu}(x)$, and the spacetime covariant derivative $\nabla_{\mu}$ acts on the spinor via vierbein and spin connection.
@@
Many extended theories of gravitation naturally host a scalar degree of freedom [26]. Regarding $\mathcal{L}_{\text{grav}}$, several possibilities exist, one being a massive Brans-Dicke action
@@
Since the matter sector in Eq. (7) has a dilatation symmetry (with $\chi$ having the length dimension $-1$), requiring the gravitational sector to also possess this symmetry would force $V(\chi)$ to take the form $\chi^{4}$. When $\omega=0$, $\mathcal{L}_{\text{grav}}$ in Eq. (11) belongs to the $f(\mathcal{R})$ family of gravity [27, 28]. It can also be embedded within a broader family of classically scale invariant gravity (or “agravity”) which includes the Weyl term, $\mathcal{C}^{\mu\nu\lambda\rho}\,\mathcal{C}_{\mu\nu\lambda\rho}$ [30, 29]. Alternatively, it can represent Horndeski gravity, the most general theory in four dimensions constructed out of the metric tensor and a scalar field that leads to second-order equations of motion [31]. Another viable option would be Lyra geometry [32], where the inherent scale function of this geometry can play the role of the scalon field $\chi$. However, the exact details of the gravitational sector are not the focus of our paper. Our work concerns the mattersector, and we require only the existence of a ‘scalon’ singlet $\chi$ that couples with the Higgs field $\Phi$ as described in Eq. (7).
While several choices for $\mathcal{L}_{\text{grav}}$ are available, we shall adopt the pure $\mathcal{R}^{2}$ theory, viz. $\omega=0$ and $V(\chi)\simeq\chi^{4}$, for the gravitational sector in the rest of this paper
$\mathcal{L}_{\text{R2}}=\chi^{2}\,\mathcal{R}-f_{0}\,\chi^{4}\,.$ (12)
At the classical level, this yields $\mathcal{R}=2f_{0}\,\chi^{2}$. Being adimensional, it has dilatation invariance, a crucial feature that proves very convenient for addressing issues in late-time cosmology, as demonstrated in this paper and its companion [14].  [^5] For pure $\mathcal{R}^{2}$ gravity, the Newtonian limit has been established [33, 34]. The theory is also known to be Ostrogradsky stable and free of ghosts [34].
A new mechanism to generate variable $\hbar$ and $c$—As a component of the gravitational sector, the scalon $\chi$ serves as a background field for the matter sector. At a given point $x^{*}$ on the manifold, its value is $\chi_{*}:=\chi(x^{*})$. In the open set vicinity of the point $x^{*}$, the terms $\frac{\mu^{2}}{2}\,\chi_{*}^{2}\,\Phi^{2}-\frac{\lambda}{4}\,\Phi^{4}$ in Eq. (7), evaluated for the prevailing $\chi_{*}$, induce a spontaneous breaking of the discrete $\mathbb{Z}_{2}$ symmetry of the Higgs singlet, given by Eq. (10)  [^6] Note that the $U(1)$ gauge symmetry described by Eqs. (8)–(9) remains unbroken in this process..This process, traditionally known as the Higgs mechanism [35], results in a non-zero Higgs vacuum expectation value (VEV)

### v2

# A mechanism to generate varying speed of light via Higgs-dilaton coupling:Theory and cosmological applications

@@ Abstract
We consider a Higgs–dilaton action which allows the Higgs field $\Phi$ to interact with a dilaton field $\chi$ of the background spacetime through the term $\chi^{2}\,\Phi^{\dagger}\Phi$. Upon spontaneous gauge symmetry breaking, the vacuum expectation value (VEV) of the Higgs field becomes proportional to $\chi$. While traditionally this linkage is employed to make the Planck mass and particle masses dependent on $\chi$, we present an alternative mechanism: the Higgs VEV will be used to construct Planck’s quantum of action $\hbar$ and speed of light $c$. Specifically, each open set vicinity of a given point $x^{*}$ on the spacetime manifold is equipped with a replica of the Glashow–Weinberg–Salam action operating with its own effective values of $\hbar_{*}$ and $c_{*}$ per $\hbar_{*}\propto\chi^{-1/2}(x^{*})$ and $c_{*}\propto\chi^{1/2}(x^{*})$, causing these “fundamental constants” to vary alongside the dynamical field $\chi$. Moreover, in each open set around $x^{*}$, the prevailing value $\chi(x^{*})$ determines the length and time scales for physical processes occurring in this region as $l\propto\chi^{-1}(x^{*})$ and $\tau\propto\chi^{-3/2}(x^{*})$. This leads to an anisotropic relation $\tau^{-1}\propto l^{-3/2}$ between the rate of clocks and the length of rods, resulting in a distinct set of novel physical phenomena. For late-time cosmology, the variation of $c$ along the trajectory of light waves from distant supernovae towards the Earth-based observer necessitates modifications to the Lemaître redshift relation and the Hubble law. These modifications are capable of: (1) Accounting for the Pantheon Catalog of Type Iasupernovae through a declining speed of light in an expanding Einstein–de Sitter universe, thus avoiding the need for dark energy; (2) Revitalizing Blanchard–Douspis–Rowan-Robinson–Sarkar’s CMB power spectrum analysis that bypassed dark energy [A&A 412, 35 (2003)]; and (3) Resolving the $H_{0}$ tension without requiring a dynamical dark energy component.
I Motivation
In light of several outstanding issues and tensions in cosmology, growing interest has emerged in promoting the parameters of physical models to be variable scalar fields in spacetime [1, 3, 2]. For example, the cosmological constant $\Lambda$ is equipped with a kinetic term that allows it to evolve and ‘relax’ to its (small) value in the current epoch, thus potentially offering a relief to the fine-tuning and coincidence problems in late-time cosmology [5, 4].
@@ I Motivation
where $\mathcal{R}$ is the Ricci scalar, and $\omega$ a dimensionless (BD) parameter. The BD theory has grown into a family of scalar–tensor theories, a very popular theme of research nowadays. The field $\chi$, often referred to as a ‘dilaton’, also naturally arises in string theory and other contexts [10].
Comparing Eq. (2) with the Einstein–Hilbert (EH) action
$\mathcal{S}_{\text{EH}}=\int d^{4}x\,\frac{\sqrt{-g}}{16\pi}\frac{c^{3}}{\hbar\,G}\,\mathcal{R}$ (3)
where the speed of light $c$ and quantum of action $\hbar$ are explicitly restored, the (Newton) gravitational ‘constant’ in BD theory becomes a scalar field via the following formal identification
$\chi^{2}:=\frac{c^{3}}{\hbar\,G}\,.$ (4)
Under the canonical assumption that $c$ and $\hbar$ do not depend on $\chi$, Eqs. (1) and (4) lead to the relationships
$G=\frac{c^{3}}{\hbar}\chi^{-2};\ \ \ M_{P}=\frac{\hbar}{c}\chi;\ \ \ l_{P}=\chi^{-1};\ \ \ \tau_{P}=\frac{1}{c}\chi^{-1}$ (5)
The dilaton field $\chi$ hence directly determines the Planck length.
It is imperative to stress that, in BD theory, the field $\chi$ itself serves as a (dynamical) length scale that is used to construct $G$, a dimensionful quantity. However, the identity expressed in Eq. (4) permits an alternative mechanism: The field $\chi$, a quantity equal to $\left(c^{3}/(\hbar G)\right)^{1/2}$, can also be used to construct $c$ and $\hbar$, rather than $G$. The key objective of our paper is to establish the relations of $\hbar$ and $c$ with $\chi$ via the matter sector.
In the original BD theory, the matter sector coupled minimally to the gravitational sector to preserve Einstein’s Equivalence Principle [9]. However, around the same time, Utiyama and DeWitt showed that even if one starts with a minimal coupling action between matter and gravity, radiative corrections would generate non-minimal coupling between them [11]. Since the 1980s, efforts to introduce non-minimal coupling directly at tree level have been actively pursued in the literature. For example, in [12, 13, 14, 15, 16] Wetterich embedded the BD action in a broader framework that projects a dilatation symmetry, with the non-minimal coupling in the form  [^1] Wetterich referred the field $\chi$ a ‘cosmon’ [12]. In place of Eq. (6), he considered a more general form $V(\chi,\Phi)=\Phi^{4}\,w\left(\chi^{2}/\Phi^{2}\right)$ where $w$ is a function of the ratio $x:=\chi^{2}/\Phi^{2}$. If $w(x)$ adopts an affine form $\lambda-\mu^{2}\,x$, then Eq. (6) would ensue.
@@
where $\lambda\in\mathbb{R^{+}}$ and $\mu\in\mathbb{R}$ are dimensionless parameters. In [17], a similar form was also considered in the context of Higgs–dilaton cosmology. Preceding to this, in [18], in place of $\chi^{2}$, Fujii let the Higgs doublet couple with the Ricci scalar in the form $\mathcal{R}\,\Phi^{\dagger}\Phi$.
@@
In the presence of Higgs–dilaton coupling, when the electroweak gauge symmetry of the action undergoes spontaneous breaking, the field $\chi$ influences the vacuum expectation value (VEV) of the Higgs doublet $\Phi$, which in turn affects the mass of the fermions, the gauge vector bosons, and the Higgs bosons in the GWS model. Specifically, Fujii and Wetterich allowed the particle mass $m$ to scale proportional to $\chi$, thereby maintaining a constant ratio $m/M_{P}$ at classical level. We shall refer to their practice collectively as the Fujii–Wetterich (FW) scheme. Moreover, the length scale $l$ and the time scale $\tau$ of any given physical process also become dependent on $\chi$ as $l\propto\chi^{-1}$ and $\tau\propto\chi^{-1}$, in exact proportion to $l_{P}$ and $\tau_{P}$ respectively. Importantly, the field $\chi$ affects not only the Planck time per Eq. (5) but also the rate of physical clocks (viz. the revolution rate of a mechanical clock, the oscillation rate of an atomic clock, or the decay rate of an unstable quantum system) which are made of matter governed by the GWS model. Quantitatively, the FW scheme thus predicts that the evolution rate $\tau^{-1}$ of physical processes is proportional to $\chi$.  [^2] We find it illuminating to quote Fujii [18]: “… the time and length in the microscopic unit frame are measured in units of $m^{-1}(t)$, in agreement with the physical situation that the time scale of atomic clocks, for example, is provided by the atomic levels which are determined by the Rydberg constant $(me^{4})^{-1}$ and Wetterich [14]: “The clock provided by the Hubble expansion in the standard description is now replaced by a clock associated to the increasing value of $\chi$.
@@
As stated earlier, our approach constitutes a major departure from the FW scheme: Instead of imposing that $\hbar$ and $c$ be constant, we require the charge and inertial mass of particles to be independent of $\chi$. This unambiguously leads to the relations $c\propto\chi^{1/2}$ and $\hbar\propto\chi^{-1/2}$, while $G$ remains constant. Furthermore, via the time evolution of quantum states $i\hbar\frac{\partial}{\partial t}\left|\psi\right\rangle=\hat{H}\left|\psi\right\rangle$, the dependency $\hbar\propto\chi^{-1/2}$, along with $\hat{H}\propto\chi$, causes the evolution rate of physical processes to scale as $\tau^{-1}\propto\chi^{3/2}$, in decisive distinction from the FW scheme (which posits $\tau^{-1}\propto\chi$).
@@
The anisotropic time scaling enabled in our approach, characterized by its anomalous $3/2$–exponent, leads to new physics with a distinct set of phenomenology andpredictions. We will discuss these significant implications in this paper, with further details developed in Ref. [19].
In brief:  History of variable $c$ and $\hbar$
@@
Prior to the development of Brans-Dicke theory, Dicke also briefly considered variability in the speed of light $c$ in 1957 [20] although he did not pursue it further. A relatively little-known fact is that the concept of variable speed of light (VSL) was initially explored by Einstein in 1911, published in three obscure papers (originally in German) [21, 22, 23]  [^3] It seems that Dicke in 1957 [20] was not aware of Einstein’s 1911 VSL proposal [21]. during his quest for a generally covariant theory of gravity. In [22, 23] Einstein emphasized that the Michelson–Morley experimental results and Lorentz symmetry are meant to hold only locally; while $c$ is invariant with respect to local boosts, its value needs not be universal. Einstein envisioned the possibility that the gravitational potential affected the clock rate, hence generating a varying $c$ in spacetime. Such a variation in $c$ could then produce a refraction effect on light rays around the Sun’s disc as a result of Huygens’s principle [21]. The later successes of General Relativity (GR) quickly overshadowed Einstein’s 1911 paper on VSL, leading him to neglect the idea, however. In the 1990s, the VSL concept was revived independently by Moffat and by Albrecht and Magueijo to tackle issues in early-time cosmology, such as the horizon paradox [24, 25]. Since then, several researchers actively explore various aspects of VSL [53, 54, 29, 26, 47, 50, 49, 48, 37, 38, 67, 51, 52, 65, 66, 81, 95, 85, 63, 62, 61, 79, 55, 30, 76, 32, 87, 91, 69, 88, 64, 45, 34, 84, 92, 89, 93, 73, 80, 82, 86, 57, 56, 70, 90, 94, 58, 71, 68, 72, 96, 46, 44, 41, 83, 78, 33, 40, 39, 36, 35, 75, 74, 42, 60, 59, 77, 97, 98, 99, 100, 43, 31, 27, 28].
@@
The possibility of a varying quantum of action $\hbar$ has been much less explored, with a rare exception of [101]. One important empirical guidance is that, although the fine-structure constant $\alpha:=e^{2}/(\hbar c)$ is known to ‘run’ in the renormalization group (RG) flow of quantum field theory, there is no clear experimental evidence supporting $\alpha$ varying in spacetime  [^4] In [102, 103] Webb et al. reported observational evidence of a varying $\alpha$ with respect to the redshift, but this result has been in contention; e.g., see a recent paper [104]. In addition, it violates the tight theoretical bound given in [105].. If $\alpha$ is to remain independent of the field $\chi$ while $c$ is allowed to vary, it would require a co-variation of $e$, $\hbar$, and $c$. One such scenario, as briefly mentioned earlier, posits $\hbar\propto\chi^{-1/2}$ and $c\propto\chi^{1/2}$, which would insulate $\alpha$ from depending on $\chi$.
II Higgs–dilaton coupling
Consider the full action in a 4-dimentional spacetime:
@@ II Higgs–dilaton coupling
Let us first focus on the matter sector. As a prototype, we consider the following Lagrangian, representing a massless spinor field $\psi$ coupled with a massless $U(1)$ gauge field $A_{\mu}$, and a charge-neutral Higgs singlet $\Phi$ coupled with a dilaton $\chi$:
@@
All parameters $\alpha$, $f$, $\mu$, and $\lambda$ are dimensionless. It is important to remark that the adimensional action of the matter sector, $\int d^{4}x\sqrt{-g}\,\mathcal{L}_{\text{mat}}$, does not involve the (dimensionful) quantum of action $\hbar$ and speed of light $c$ at its outset. The length dimensions of $\{\psi,\,A_{\mu},\,\Phi$} are $\{-3/2,-1,-1\}$ respectively. The gamma matrices satisfy $\gamma^{\mu}\gamma^{\nu}+\gamma^{\nu}\gamma^{\mu}=2g^{\mu\nu}$, and the spacetime covariant derivative $\nabla_{\mu}$ acts on the spinor via vierbein and spin connection.
@@
Regarding the gravitation Lagrangian $\mathcal{L}_{\text{grav}}$, most extended theories of gravitation naturally host a scalar degree of freedom [106, 107, 108, 10, 109, 110]. Several possibilities exist, one being a massive Brans-Dicke action
@@
When $\omega=0$, $\mathcal{L}_{\text{grav}}$ in Eq. (12) belongs to the $f(\mathcal{R})$ family of gravity [111, 112, 113]. It can also be embedded within a broader family of scale invariant gravity, such as those considered in [114, 115, 116] or “agravity” which includes the Weyl term, $\mathcal{C}^{\mu\nu\lambda\rho}\,\mathcal{C}_{\mu\nu\lambda\rho}$ [118, 117, 119]. Alternatively, it can represent Horndeski gravity, the most general theory in four dimensions constructed out of the metric tensor and a scalar field that leads to second-order equations of motion [120]. Another viable option would be Lyra geometry [121], where the inherent scale function of this geometry can play the role of the dilaton field $\chi$. However, the exact details of the gravitational sector are not the focus of our paper. Our work concerns the matter sector, and we require only the existence of a ‘dilaton’ singlet $\chi$ that couples with the Higgs field $\Phi$ as described in Eq. (8).
It has been established in [122, 114] that a scale-invariant action, such as the one described in Eqs. (7), (8), and (12), can evade observational constraints on the fifth force. Among the several possible choices for $\mathcal{L}_{\text{grav}}$, the pure $\mathcal{R}^{2}$ theory $\mathcal{L}_{\text{R2}}=\chi^{2}\,\mathcal{R}-f_{0}\,\chi^{4}$ is a promising candidate [123, 124, 125, 126]. At the classical level, the dilaton and the Ricci scalar are one-to-one related as $\mathcal{R}=2f_{0}\,\chi^{2}$. The Newtonian limit has been established [127, 123]. The theory is known to be Ostrogradsky stable and free of ghosts [123], and properties of its graviton propagators have also been investigated [123, 128, 124].
III A new mechanism to generate variable $\hbar$ and $c$
Assuming that $\chi$ is a slowly varying background field, at a given point $x^{*}$ on the manifold, its value is $\chi_{*}:=\chi(x^{*})$. In the open set vicinity of the point $x^{*}$, the terms $\frac{\mu^{2}}{2}\,\chi_{*}^{2}\,\Phi^{2}-\frac{\lambda}{4}\,\Phi^{4}$ in Eq. (8), evaluated for the prevailing $\chi_{*}$, induce a spontaneous breaking of the discrete $\mathbb{Z}_{2}$ symmetry of the Higgs singlet, given by Eq. (11)  [^5] Note: The $U(1)$ gauge symmetry described by Eqs. (9)–(10) remains unbroken in this process..This process, traditionally known as the Higgs mechanism [129, 131, 132, 130], results in a non-zero vacuum expectation value (VEV) of the Higgs field